# Unsupervised Video Highlight Detection by Learning from Audio and Visual Recurrence

### v1

@@ Abstract
With the exponential growth of video content, the need for automated video highlight detection to extract key moments or highlights from lengthy videos has become increasingly pressing. This technology has the potential to significantly enhance user experiences by allowing quick access to relevant content across diverse domains. Existing methods typically rely either on expensive manually labeled frame-level annotations, or on a large external dataset of videos for weak supervision through category information. To overcome this, we focus on unsupervised video highlight detection, eliminating the need for manual annotations. We propose an innovative unsupervised approach which capitalizes on the premise that significant moments tend to recur across multiple videos of the similar category in both audio and visual modalities. Surprisingly, audio remains under-explored, especially in unsupervised algorithms, despite its potential to detect key moments. Through a clustering technique, we identify pseudo-categories of videos and compute audio pseudo-highlight scores for each video by measuring the similarities of audio features among audio clips of all the videos within each pseudo-category. Similarly, we also compute visual pseudo-highlight scores for each video using visual features. Subsequently, we combine audio and visual pseudo-highlights to create the audio-visual pseudo ground-truth highlight of each video for training an audio-visual highlight detection network. Extensive experiments and ablation studies on three highlight detection benchmarks showcase the superior performance of our method over prior work.
@@ 2 Related Work
Most of the existing approaches for highlight detection rely on manually annotated frame-level supervision [12, 15, 34, 43, 47, 1, 23]. However, these annotations are laborious and expensive to obtain. To address this limitation, some works focus on using only video-level tags or category information as weak supervision [45, 3, 13, 29, 44, 46]. Several works adopt ranking frameworks. For example, LM [44] exploits video duration and ranks clips from shorter videos higher, whereas, MN [13] ranks clips belonging to a given category higher than those from other categories. Ye et al. [46] leverages temporal reasoning and encodes cross-modal relationships using an efficient audio-visual tensor fusion mechanism. However, these methods require training on large-scale external data. There is a recent unsupervised method based on contrastive learning [2] which does not utilize any external data. However, they do not exploit audio cues for unsupervised learning. Some recent works explore supervised query-based highlight detection, where the goal is to extract highlight relevant to a given textual query [21, 23, 27]. In contrast, we focus on unsupervised audio-visual highlight detection which does not rely on any text input.
@@ 3 Our Approach
Let’s say we have a set of $M$ unlabeled videos $\{V_{j}\}_{j=1}^{M}$. We split each video $V_{j}$ into clips with an equal number of frames, resulting in $n$ clips. From each clip $c_{i}$ (where $i=1,2,...n$), we extract its corresponding visual features $v_{i}\in\mathbb{R}^{d_{v}}$ using a pre-trained visual feature extractor and audio features $a_{i}\in\mathbb{R}^{d_{a}}$ using a pre-trained audio feature extractor. Each video $V_{j}$ can then be represented using its corresponding set of visual features, $\{v_{i}\}_{i=1}^{n}$, and audio features, $\{a_{i}\}_{i=1}^{n}$. The aim of our highlight detection method is to predict a set of highlight scores $\{h_{i}\}_{i=1}^{n}$, where $h_{i}$ indicates the highlight score of each clip $c_{i}$ in video $V_{j}$.
In the following, we discuss the three broad steps mentioned above in details.
@@ 3.2 Generating Audio-Visual Pseudo-Highlight
We generate an audio-visual pseudo-highlight for each video, which we use to train our model. For each video, we first compare each of its clips with all the clips across videos of the same pseudo-category using their audio and visual features to obtain audio pseudo-highlight and visual pseudo-highlight scores, respectively, of the video. We then aggregate these scores to obtain the audio-visual pseudo-highlight scores. Finally, we select the clips with the highest audio-visual pseudo-highlight scores to generate the audio-visual pseudo-highlight of the video.
Audio Pseudo-Highlight (A-PH): Let’s consider there are $K$ videos assigned to a particular pseudo-category, and $\{a_{k}\}_{k=1}^{S}$ represents a set of audio features corresponding to all the $S$ clips across all $K$ videos of that pseudo-category. We assign an audio pseudo-highlight score $aph_{i}$ to the $i$-th clip of a video based on how repetitive or similar the corresponding audio features $a_{i}$ are in the videos of that pseudo-category (referred to as audio recurrence). To achieve this, we utilize cosine similarity to compare the audio features of a clip $a_{i}$ in a video with all the clips belonging to the pseudo-category. We can write $aph_{i}$ computation as follows:
@@ 3.3 Audio-Visual Highlight Detection Network
$\{v_{i}^{v}\}_{i=1}^{n}=\text{Attn}_{v\rightarrow v}(\{v_{i}\}_{i=1}^{n})$ (3)
@@
We optimize our network by minimizing the binary cross-entropy loss between the predicted highlight scores and the audio-visual pseudo-highlight (AV-PH).
@@ 4.1 Datasets and Settings
We evaluate our proposed method using three benchmark highlight detection datasets: YouTube Highlights [41], TVSum [40], and QVHighlights [21]. YouTube Highlights is constructed by mining YouTube videos related to six specific categories - dog, parkour, gymnastics, skiing, skating, and surfing with about 100 videos in each category. We utilize the train and test splits provided in this dataset. TVSum has 50 videos across 10 diverse categories such as changing vehicle tire, making sandwich, and so on. Following prior works [1, 2, 35], we randomly split this dataset with $80\%$ of the videos for training and $20\%$ for testing, and we run our experiments on this dataset five times and report the average performance. QVHighlights is much larger than the other two datasets, with over 10,000 videos. It is primarily designed for query-focused video highlight detection and moment retrieval. Each video is associated with a textual query and corresponding saliency/highlight scores. The dataset comes with standard train, validation, and test splits with a ratio of 70:15:15. Since our method only requires videos, we ignore the user query annotations. For a fair comparison, on QVHighlights, we evaluate our method against prior non-query-based methods.
@@
Evaluation Metrics: Following prior works on QVHighlights [23, 21], we report our performance using Mean Average Precision (mAP), which considers the highlight scores for all the clips, and HIT@1, which considers the hit ratio of the clip with the highest score for each video. We consider only the clips rated as Very Good by users for evaluation. On YouTube, we evaluate using mAP, and on TVSum, we report the mean average precision on the top five predicted clips (top-5 mAP) as in prior works [2, 1]. All metrics are reported as percentages.
Implementation Details: We implement our models using PyTorch [31]. For YouTube and TVSum, we utilize one self-attention and one bi-modal attention module in our audio-visual highlight detector network, as in the previous method [1]. However, since QVHighlights is a much larger dataset, following previous studies [23], we introduce an additional self-attention module and fully connected layer in the score regressor to handle the increased complexity and scale of the data. We use Adam to optimize [18] our models. We train our models for 20 epochs with a learning rate of $2.5\times 10^{-3}$ on TVSum and for 100 epochs with a learning rate of $5\times 10^{-3}$ on YouTube. For QVHighlights, we train our models with a learning rate of $5\times 10^{-4}$ for $10$ epochs. Following prior works [2, 13], we select the top $t=50\%$ clips based on the audio-visual pseudo-highlight scores (AV-PH) to create pseudo-highlights for training.
@@ 4.2 Highlight Detection Results
Table 3: We compare with prior methods that utilize external data. Even without the advantage of external data, our method outperforms three of these methods.
@@ 4.3 Ablation Studies
We conduct extensive ablation studies to analyze the relative impact and contribution of each component of our approach. For QVHighlights, we adhere to prior work [21, 23] and utilize its val split due to restrictions on the number of submissions to the evaluation server. For TVSum and YouTube, we use the test split. The following baselines are defined for our ablation experiments.
@@
Comparison with Supervised Setting: In Table 9, we compute the highlight detection performance when real ground-truths are used for training our model instead of the audio-visual pseudo-highlights. This denotes the fully supervised version of our model. Interestingly, our unsupervised method demonstrates strong performance compared to its supervised counterpart. Specifically, on YouTube, our method underperforms by only 2% compared to the supervised model. This showcases the effectiveness of the proposed pseudo-highlight mechanism as an alternative supervisory signal for training a highlight detection model if manual annotations are not available.
@@
How good are audio-visual pseudo-highlight as supervision? We conduct an experiment where we directly use the audio-visual pseudo-highlight scores of each video as predictions and compare them with predictions of our highlight detection model trained using audio-visual pseudo-highlights. Concretely, we first assign each test video to its corresponding pseudo-category by finding the nearest cluster centroid in the clustering model learned on train data in Sec. 3.1. Then, we obtain the audio-visual pseudo-highlight scores of each test video by comparing them with all training videos of the same pseudo-category. We directly use audio-visual pseudo-highlight scores of a test video as its highlight predictions. In Table 12, we compare these highlight predictions with highlights predicted by our trained model. Despite potential noise in audio-visual pseudo-highlight supervision, the highlight detection model is able to learn from it during training, indicating its worth and benefits, and achieves superior performance compared to using audio-visual pseudo-highlight directly as predictions.
@@ 5 Conclusion
We introduce a novel unsupervised audio-visual highlight detection framework. Specifically, our core idea is based on the premise that videos of similar categories tend to contain key moments that are repetitive in both audio and visual modalities across multiple videos. Leveraging this observation, we construct audio-visual pseudo-highlights to train our model. Extensive experiments and ablation studies showcase the effectiveness of our framework. Moreover, experiments also reveal that cues from the often overlooked yet informative audio modality, when coupled with the visual modality, lead to a significant improvement in unsupervised learning of video highlight detection.

### v2

@@ Abstract
With the exponential growth of video content, the need for automated video highlight detection to extract key moments or highlights from lengthy videos has become increasingly pressing. This technology has the potential to enhance user experiences by allowing quick access to relevant content across diverse domains. Existing methods typically rely either on expensive manually labeled frame-level annotations, or on a large external dataset of videos for weak supervision through category information. To overcome this, we focus on unsupervised video highlight detection, eliminating the need for manual annotations. We propose a novel unsupervised approach which capitalizes on the premise that significant moments tend to recur across multiple videos of the similar category in both audio and visual modalities. Surprisingly, audio remains under-explored, especially in unsupervised algorithms, despite its potential to detect key moments. Through a clustering technique, we identify pseudo-categories of videos and compute audio pseudo-highlight scores for each video by measuring the similarities of audio features among audio clips of all the videos within each pseudo-category. Similarly, we also compute visual pseudo-highlight scores for each video using visual features. Then, we combine audio and visual pseudo-highlights to create the audio-visual pseudo ground-truth highlight of each video for training an audio-visual highlight detection network. Extensive experiments and ablation studies on three benchmarks showcase the superior performance of our method over prior work.
@@ 2 Related Work
Most of the existing approaches for highlight detection rely on manually annotated frame-level supervision [12, 15, 34, 43, 47, 1, 23]. However, these annotations are laborious and expensive to obtain. To address this limitation, some works focus on using only video-level tags or category information as weak supervision [45, 3, 13, 29, 44, 46]. Several works adopt ranking frameworks. For example, LM [44] exploits video duration and ranks clips from shorter videos higher, whereas, Ye et al. [46] leverages temporal reasoning and encodes cross-modal relationships using an efficient audio-visual tensor fusion mechanism. However, these methods require training on large-scale external data. A recent unsupervised method based on contrastive learning [2] does not utilize any external data. However, they do not exploit audio for unsupervised learning. Some recent works explore supervised query-based highlight detection, where the goal is to extract highlight relevant to a given textual query [21, 23, 27]. In contrast, we focus on unsupervised audio-visual highlight detection which does not rely on any text input.
@@ 3 Our Approach
Let’s say we have a set of $M$ unlabeled videos $\{V_{j}\}_{j=1}^{M}$. We split each video $V_{j}$ into clips with an equal number of frames, resulting in $n$ clips. From each clip $c_{i}$ (where $i=1,2,...n$), we extract its corresponding visual features $v_{i}\in\mathbb{R}^{d_{v}}$ using a pre-trained visual feature extractor and audio features $a_{i}\in\mathbb{R}^{d_{a}}$ using a pre-trained audio feature extractor. Each video $V_{j}$ can then be represented using its corresponding set of visual features, $\{v_{i}\}_{i=1}^{n}$, and audio features, $\{a_{i}\}_{i=1}^{n}$. The aim of our highlight detection method is to predict a set of highlight scores $\{h_{i}\}_{i=1}^{n}$, where $h_{i}$ indicates the highlight score of each clip $c_{i}$ in video $V_{j}$. In the following, we discuss the three broad steps in our approach.
@@ 3.2 Generating Audio-Visual Pseudo-Highlight
We generate an audio-visual pseudo-highlight for each video to train our model. For each video, we first compare each of its clips with all the clips across videos of the same pseudo-category using their audio and visual features to obtain audio pseudo-highlight and visual pseudo-highlight scores, respectively. We then aggregate these scores to generate audio-visual pseudo-highlight of the video.
Audio Pseudo-Highlight (A-PH): Let, there are $K$ videos assigned to a particular pseudo-category, and $\{a_{k}\}_{k=1}^{S}$ represents a set of audio features corresponding to all the $S$ clips across all $K$ videos of that pseudo-category. We assign an audio pseudo-highlight score $aph_{i}$ to the $i$-th clip of a video based on how repetitive the corresponding audio features $a_{i}$ are in the videos of that pseudo-category (referred to as audio recurrence). To achieve this, we utilize cosine similarity to compare the audio features of a clip $a_{i}$ in a video with all the clips in the pseudo-category. We can write $aph_{i}$ computation as follows:
@@ 3.3 Audio-Visual Highlight Detection Network
$\small\{v_{i}^{v}\}_{i=1}^{n}=\text{Attn}_{v\rightarrow v}(\{v_{i}\}_{i=1}^{n})$ (3)
@@
The training procedure for our network follows the standard approach used in the supervised network [1]. However, instead of using ground-truth annotations, we use audio-visual pseudo-highlights as the supervisory signal to train our network. We optimize our network by minimizing the binary cross-entropy loss between the predicted highlight scores and the audio-visual pseudo-highlights (AV-PH).
@@ 4.1 Datasets and Settings
We evaluate our method using three benchmark highlight detection datasets: YouTube Highlights [41], TVSum [40], and QVHighlights [21]. YouTube Highlights is constructed by mining YouTube videos related to six specific categories such as parkour, gymnastics, skiing, and so on, with about 100 videos in each category. We utilize the train and test splits provided in this dataset. TVSum has 50 videos across 10 diverse categories. Following prior works [1, 2, 35], we randomly split this dataset with $80\%$ of the videos for training and $20\%$ for testing, and we run our experiments on this dataset five times and report the average performance. QVHighlights is larger with over 10,000 videos. It is primarily designed for query-focused video highlight detection and moment retrieval. Each video is associated with a textual query and corresponding saliency/highlight scores. The dataset comes with standard train, validation, and test splits with a ratio of 70:15:15. Since our method only requires videos, we ignore the user query annotations. For a fair comparison, on QVHighlights, we evaluate our method against prior non-query-based methods.
@@
Evaluation Metrics: Following prior works on QVHighlights [23, 21], we report our performance using Mean Average Precision (mAP), which considers the highlight scores for all the clips, and HIT@1, which considers the hit ratio of the clip with the highest score for each video. We consider only the clips rated as Very Good by users for evaluation. On YouTube, we evaluate using mAP, and on TVSum, we report mAP on the top five predicted clips (top-5 mAP) as in prior works [2, 1]. All metrics are reported as percentages.
Implementation Details: We implement our models using PyTorch [31]. For YouTube and TVSum, we utilize one self-attention and one bi-modal attention module in our audio-visual highlight detector network, as in the previous method [1]. However, since QVHighlights is a much larger dataset, following previous studies [23], we introduce an additional self-attention module and fully connected layer in the score regressor to handle the increased complexity and scale of the data. We use Adam to optimize [18] our models. We train our models for 20 epochs with a learning rate of $2.5\times 10^{-3}$ on TVSum and for 100 epochs with a learning rate of $5\times 10^{-3}$ on YouTube. For QVHighlights, we train our models with a learning rate of $5\times 10^{-4}$ for $10$ epochs. As mentioned in Sec. 3.1, we empirically select the number of clusters, $K$, by maximizing the Silhouette Coefficient. For all three datasets, we search in the range of 4 to 15 to find the value of $K$. For YouTube, TVSum, and QVHighlights, we find the optimal values of $K$ to be 6, 8, and 7, respectively. Following prior works [2, 13], we select the top $t=50\%$ clips based on the audio-visual pseudo-highlight scores (AV-PH) to create pseudo-highlights for training.
@@ 4.2 Highlight Detection Results
Table 3: We compare with prior methods on TVSum that utilize external data. Even without the advantage of external data, our method outperforms three of these methods.
@@ 4.3 Ablation Studies
We conduct extensive ablation studies to analyze the relative impact of each component of our approach. For QVHighlights, we follow prior work [21, 23] and utilize its val split due to limited number of submissions to the evaluation server. For TVSum and YouTube, we use the test split. We define the following baselines for our experiments.
@@
Impact of different number of clusters, $K$: In Table 9, we compare the performance of our method for different values of number of clusters, $K$. Results show a performance drop for non-optimal values of $K$, while the optimal values chosen by our method in Sec. 3.1 yield best results.
Number of clusters 6 7 8
TVSum 52.84 54.76 60.34
YouTube 68.30 63.96 63.56
QVHighlights 17.80 18.41 18.02
Table 9: Ablation on the different number of clusters, $K$. The optimal values of $K$ selected by our method yield best results.
Comparison with Supervised Setting: In Table 10, we compute the highlight detection performance when real ground-truths are used for training our model instead of the audio-visual pseudo-highlights. This denotes the fully supervised version of our model. Interestingly, our unsupervised method demonstrates strong performance compared to its supervised counterpart. On YouTube, our method underperforms by only 2% compared to the supervised model. This showcases the effectiveness of the proposed pseudo-highlight mechanism as an alternative supervisory signal for highlight detection if manual annotations are not available.
@@
Audio-visual pseudo-highlight as supervision: We conduct an experiment where we directly use the audio-visual pseudo-highlight scores of each video as predictions and compare them with predictions of our highlight detection model trained using audio-visual pseudo-highlights. Concretely, we first assign each test video to its corresponding pseudo-category by finding the nearest cluster centroid in the clustering model learned on train data in Sec. 3.1. Then, we obtain the audio-visual pseudo-highlight scores of each test video by comparing them with all training videos of the same pseudo-category. We directly use audio-visual pseudo-highlight scores of a test video as its highlight predictions. In Table 13, we compare these highlight predictions with highlights predicted by our trained model. Despite potential noise in pseudo-highlight supervision, our model is able to learn from it during training, indicating its worth and benefits, and achieves superior performance compared to using pseudo-highlight directly as predictions.
@@ 5 Conclusion
We introduce a novel unsupervised audio-visual highlight detection framework. Our core idea is based on the premise that videos of similar categories tend to contain key moments that are repetitive in both audio and visual modalities across multiple videos. Leveraging this observation, we construct audio-visual pseudo-highlights to train our model. Extensive experiments showcase the effectiveness of our framework and reveal that cues from the often overlooked yet informative audio modality, when coupled with the visual modality, lead to a significant improvement in unsupervised learning of video highlight detection.
Acknowledgements: Zahidul Islam and Mrigank Rochan acknowledge the support of the University of Saskatchewan and the Natural Sciences and Engineering Research Council of Canada (NSERC).